17 years ago I was half the age I am now. How old am I?
Answer by nyc_function(2741)

Age now = x
Age 17 years ago = x - 17

Then (1/2)x = x - 17 is the equation we need to find your current age.

Multiply both sides of the equation by the denominator 2.

After doing so, we get x = 2x - 34.

We now solve for x.

x = 2x - 34

x - 2x = -34

-x = -34

x = -34/-1

x = 34

Done!